Tohban Report 2014-04-30

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 23 Apr 2014
End Date: 30 Apr 2014
Tohban: Hazel Bain (Vice tohban: Hugh Hudson)
Tohban email:
Next Tohban: Albert Shih (Vice Tohban: Brian Dennis)
List all reports


Solar Activity

Solar activity at the start of the week was moderate with a number of C class flares. On the 25th of April AR 12046 produced an isolated X1.6 class flare, the active region was occulted at the time of the event. RHESSI caught the rise, peak and decay of the flare. Following that, solar activity has been low for the remainder of the week with only a few additional C class flares. There are currently 4 ARs on disk only one of which AR 12049 produced a C class flare in the last 24 hours

How many GOES flares occurred?

 Flares above B, C, M, X class were      8    15     0     1

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were      4    15     0     1

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were      0     1     0     1

There were RHESSI flares/GOES flares 127 / 24 over the time range 23-Apr-14 30-Apr-14

Memory Management

The SSR fill level...

Spacecraft Status

The spacecraft switched to IDLE mode on Apr 26th. This was most likely due to a spin rate anomaly when the RHESSI spin axis is parallel to the local magnetic field line. RHESSI was commanded back to precession mode at 2:17pm PST on the 26th April.

On the 28th April RHESSI flew threw an eclipse 3 times.

Several spikes were observed in the quicklook plots from ~1400UT 27th Apr - ~0600 UT 28th Apr.This is most likely due to noise introduced in the Wallops downlink. The data was not replayed since the Sun was quiet at this time.

Data Gaps

No data gaps.

Detector issues

The rear fast threshold on detector 8 was increased from 0xF0 to 0xFF (max value) on April 23, at 18:40 UTC.

The rear reset rate is starting to creep up again for detectors 4, 8 and 9.

Commands sent

The rear fast threshold on detector 8 was increased from 0xF0 to 0xFF (max value) on April 23, at 18:40 UTC.

Other notes

A second/vice tohban was introduced this week, since there are so many detector and related issues to watch out for that it's getting to be a bit much for one person to do. The second tohban will provide a second pair of eyes watching for things like spurious attenuator changes, increases in detector noise, or other problems. Since there's an extremely limited number of tohbans at SSL, the second tohban will be an outside volunteer (which also gives better time coverage).

Decimation Active/Vigorous
HLAT Decimation Rear decimation weight 6, no front decimation
Night time data (fronts) +/- 4 minutes
Night time data (rears) +/- 4 minutes
Require extra passes? No
Requirement for moving pointer? No
Attenuator operation Normal
Detector problems? 4, 8, 9
Personal tools